OpenStudy (amonoconnor):

When evaluating the integral of "(cosx)/((2 + sinx)^2)(dx)", with upper and lower lower limits of Pi/2 & -Pi/2 respectively, what is the answer and what are the steps? I know how to work these problems, generally, but am having a some hiccups on a couple questions. Any and all help is greatly appreciated! :)

OpenStudy (amonoconnor):

So far I have found: u = (2 + sinx)^2 du = 2(2 + sinx)(cosx)(dx)

OpenStudy (anonymous):

\[\int\limits_{\frac{ -\pi }{ 2 }}^{\frac{ \pi }{ 2 }}\frac{ \cos x }{\left( 2+\sin x \right)^2 }dx\]

OpenStudy (amonoconnor):

Also, cos(dx) = du/[2(2 + sinx)]

OpenStudy (amonoconnor):

Correct!

OpenStudy (praxer):

it is better if you substitute (2+sinx)=t and than find the derivative.

OpenStudy (anonymous):

put 2+sin x=t cos x dx=dt \[when~x=-\pi/2,t=2+\sin(-\pi/2) ,t=2-\sin \frac{ \pi }{2 }=2-1=1\] \[when~ x=\frac{ \pi }{2 },t=2+\sin \frac{ \pi }{2 }=2+1=3\] solve.

OpenStudy (anonymous):

\[I=\int\limits_{1}^{3} t ^{-2}~dt\]=?

OpenStudy (amonoconnor):

...when do you use the anti derivative of the integral, after u (or t) substitution, and then subtract the anti.deriv. with the lower limit plugged into it, from the anti.deriv. With the upper limit plugged into it?